OpenStudy (anonymous):

If a/d=c/d, b = 0, and d = 0, then ad=______ by the Means-Extremes Product Property. Answers. a. cd b. ab c. dc d. bc

OpenStudy (anonymous):

is it d. bc?

jimthompson5910 (jim_thompson5910):

is a/d=c/d supposed to be a/b=c/d notice how I swapped out the first 'd' with 'b'

OpenStudy (anonymous):

oops i mean't a/b = c/d I just caught on to that.

jimthompson5910 (jim_thompson5910):